OpenStudy (anonymous):

[7.02] Identify the center and the radius of the circle with equation (x - 6)2 + y2 = 49.

OpenStudy (anonymous):

The general equation of a circle is: (x - h)^2 + (y - k)^2 = r^2 The equation above fits this form as follows: (x - 6)^2 + (y - 0)^2 = 7^2 Therefore: Center (h,k) = (6,0) Radius = 7

OpenStudy (anonymous):

Thank You so much!